# HOT: An Efficient Halpern Accelerating Algorithm for Optimal Transport Problems

### v1

@@ III-A HOT: A Halpern accelerating method for solving OT problem
For ease of notation, denote $w:=(y,z,x)$. A fast Halpern accelerating method [35, 43] for solving OT problems is presented in Algorithm 2. A detailed derivation of the algorithm in its current form and more discussions can be found in [35] and the references therein.
@@
Note that Step 4 in Algorithm 2 is from the Halpern iteration with a stepsize of $\frac{1}{k+2}$. Without Step 4, the HOT algorithm reduces to the ADMM with a unit step size. The following proposition shows the global convergence of the HOT algorithm. The proof of this proposition follows from [35, Corollary 3.5].
@@
for any $w=(y,z,x)\in\mathbb{R}^{M_{3}}\times\mathbb{R}^{N}\times\mathbb{R}^{N}$. Note that $\mathcal{R}(w^{*})=0$ is equivalent to the facts that $x^{*}\in\mathbb{R}^{N}$ and $(y^{*},z^{*})\in\mathbb{R}^{M_{3}}\times\mathbb{R}^{N}$ are the solution to problems (8) and (9), respectively. The KKT residual $\|\mathcal{R}(\cdot)\|$ is a commonly used and practical measure for the quality of the approximation solution to (8). It follows from [35, Theorem 3.7] that the HOT algorithm enjoys an appealing $O(1/k)$ nonergodic convergence rate in terms of the KKT residual for solving (8), which is summarized in the following proposition.
@@
$\|\mathcal{R}(\bar{w}^{k})\|\leq\left(\frac{\sigma+1}{{\sigma}}\right)\frac{R_{0}}{(k+1)}.$ (14)
@@
Note that, without acceleration, the ADMM has an $O(1/\sqrt{k})$ non-ergodic rate in terms of both the objective function value gap and feasibility violations [44, 45]. In contrast, the HOT algorithm in Algorithm 2 achieves an $O(1/k)$ non-ergodic convergence rate, offering significant advantages for solving large-scale OT problems.
@@ III-B A fast implementation of the HOT algorithm
Therefore, the main computational bottleneck of the HOT algorithm for solving (8) is solving the linear system (15). Note that the dimension of the matrix $AA^{\top}$ is $M_{3}\times M_{3}$ with $M_{3}$ defined in (4). In applications, $M_{3}$ is usually a huge number. As an illustrative example, for an image with $256\times 256$ pixels, $M_{3}=196,607$. As a result, it is not computationally affordable for computing a (sparse) Cholesky decomposition for the matrix $AA^{\top}$ or solving the linear system (15) with standard direct solver. Indeed, it is computationally expensive even for computing the matrix $AA^{\top}$. Instead, in the remaining part of this subsection, we will derive a linear time complexity procedure for solving the linear equation $AA^{\top}{y}=R$ with a given vector $R\in\mathbb{R}^{M_{3}}$. It is worthwhile mentioning that our procedure does not require calculating nor storing the matrix $AA^{\top}$. Note that ${A}A^{\top}$ can be written in the following form:
@@
$O\left(\left(\frac{1+\sigma}{\sigma}\left(\left\|x^{0}-x^{*}\right\|+\sigma\left\|s^{0}-s^{*}\right\|\right)\right)\frac{m^{2}n+mn^{2}}{\varepsilon}\right).$
@@
Note that the complexity of the Sinkhorn method for solving the original OT problem (1) is $O\left(\frac{m^{2}n^{2}}{\varepsilon^{2}}\right)$ to achieve an $\varepsilon$-accuracy solution in terms of objective function value [13]. Even by leveraging the separable structure of the cost function $c$ defined in (2), the improved Sinkhorn method, as mentioned in [3] and [32], still requires $O\left(\frac{m^{2}n+mn^{2}}{\varepsilon^{2}}+m^{2}n^{2}\right)$ to compute the KW distance. In contrast, the HOT algorithm exhibits lower overall computational complexity, offering a significant advantage in calculating the KW distance between large-scale histograms.
@@ IV Experiments
In this section, we comprehensively compare the HOT algorithm with five other state-of-the-art methods on the popular DOTmark dataset [46]. Additionally, employing the transport plan derived from Algorithm 1, we test the performance of the reduced OT model on the color transfer task. The findings reveal that HOT provides significant advantages over state-of-the-art methods in both memory and computational efficiency, particularly for large-scale problems.
@@ IV-A Numerical comparison on the DOTmark dataset
The DOTmark dataset [46] is a comprehensive collection of benchmark instances for evaluating and comparing algorithms in the field of optimal transport. It consists of a variety of instances categorized into different classes, such as Classic Images, Shapes, and Gaussian Distributions. In this experiment, we selected eight images each from the Classic Images and Shapes categories, as illustrated in Fig.2. These selected images were resized to five different resolutions: $64\times 64$, $128\times 128$, $256\times 256$, and $512\times 512$. Finally, we randomly selected 10 pairs from each category and computed the KW distance for each pair.
@@
  • Sinkhorn [10] is a widely used algorithm for computing an approximate KW distance by solving an entropy-regularized OT problem. As the Introduction Section discusses, its performance is sensitive to the regularization parameter, denoted as $\lambda$. To achieve varying levels of solution accuracy, we selected the $\lambda$ to be $1\%$, $0.1\%$, and $0.01\%$ of the median transport cost, following the setup from [3]. Due to the potential numerical instability caused by small $\lambda$ values, we employed the Log-domain Sinkhorn algorithm implemented by POT [47] as the baseline method.
  • Improved Sinkhorn [3, 33] is an improved version of the Sinkhorn algorithm, optimized for computing distances over regular two-dimensional grids. Specifically, since the kernel matrix used by the Sinkhorn algorithm can be constructed using a Kronecker product of two smaller matrices, a matrix-vector product using a matrix of dimension $M\times M$ can be replaced by two matrix-matrix products over matrices of dimension $\sqrt{M}\times\sqrt{M}$, resulting in a significant improvement in computational efficiency. In our experiment, based on the MATLAB implementation of the improved Sinkhorn in [32], we further developed the Log-domain improved Sinkhorn using PyTorch. The regularization parameter $\lambda$ is kept the same as that of the original Sinkhorn method.
@@
  • Network Simplex implemented in the Lemon C++ graph library  [^2] https://lemon.cs.elte.hu/, is a highly efficient algorithm for solving uncapacitated minimum cost flow problems [48]. It has demonstrated the computational advantages over the Sinkhorn type methods in [32] for small size images by solving the reduced model (8).
  • ADMM [41, 40] is a popular first-order primal-dual method for solving large-scale optimization problems. It has shown great potential in solving large-scale optimal transport problems (1) in a GPU setting [20]. In this numerical experiment, we use a generalized ADMM [49, 50] to solve the equivalent model (8) (replacing Step 4 in Algorithm 2 with $w^{k+1}=(1-\rho)w^{k}+\rho{\bar{w}}^{k}$ and setting $\rho=1.7$) as the baseline to evaluate the acceleration effect of the Halpern iteration.
@@
${\rm KKT_{res}}=\max\left\{\frac{\|A^{\top}y+z-c\|}{1+\|c\|},\frac{\|\min(x,z)\|}{1+\|x\|+\|z\|},\frac{\|Ax-b\|}{1+\|b\|}\right\}.$ (29)
@@
We present the average results of 10 pairs within each category for all tested algorithms in Table II. Since Gurobi runs out of memory for images sized from $256\times 256$ to $512\times 512$, we only report the feaserr for the HOT, Network Simplex, and ADMM. Additionally, Table II only shows the results of Sinkhorn-type methods with $\lambda=0.01\%$ of the median transport cost, as this parameter returns a solution of comparable quality to other methods. For more results of the improved Sinkhorn with different regularization parameters, refer to Table III. Due to space constraints, we omit the results of the Sinkhorn with varying parameters, which exhibit similar performance to the improved Sinkhorn in terms of the gap.
@@
    Computational efficiency: HOT can return a comparable solution in terms of the gap and feaserr in the shortest time. Although the Network Simplex and Gurobi have computational advantages for computing the KW distance for small-scale images, these methods cannot handle large-scale problems effectively. IPM implemented in Gurobi requires solving a linear equation that relies heavily on Cholesky decomposition, causing the computational cost of each iteration to increase rapidly with image size. Additionally, the inherent sequential nature of the Network Simplex algorithm makes it challenging to parallelize effectively. In contrast, HOT and improved Sinkhorn benefit from low per-iteration costs and are easily parallelizable. However, the improved Sinkhorn needs to recover the solution to the original problem (1) to compute the KW distance, making it unsuitable for large-scale problems. As a result, for images sized $128\times 128$, HOT achieves a 17.44x speedup over Network Simplex, a 15.83x speedup over Gurobi, and a 19.54x speedup over improved Sinkhorn. Additionally, compared to ADMM, HOT benefits from a superior $O(1/k)$ iteration complexity and saves 40% of iterations for images sized $512\times 512$.
@@
TABLE III: The numerical results of the improved Sinkhorn method with different $\lambda$.
@@
Improved Sinkhorn Classic $64\times 64$ time(s) 16.18 1.60 0.19
@@
To further illustrate the benefit of the explicit solution of the linear system presented in Proposition 4, we conduct a comparison between solving the linear equation using Proposition 4 and sparse Cholesky decomposition. Given that the average iteration number of HOT is around 1500, we solve the linear system (17) 1500 times for different image sizes. The results are shown in Fig.3. It is clear that the computational costs of Cholesky decomposition increase rapidly as image size increases. Additionally, the forward-backward substitution method used to solve the linear system is also time-consuming because it cannot be efficiently parallelized. In contrast, solving linear systems using Proposition 4 can be easily parallelized. It consumes significantly less time and remains constant as the size varies because it only requires $O(M_{3})$ flops, as shown in Corollary 1.

### v2

@@ III-A HOT: A Halpern accelerating method for solving OT problem
For ease of notation, denote $w:=(y,z,x)$. A fast Halpern accelerating method [sun2024accelerating, halpern1967fixed] for solving OT problems is presented in Algorithm 2. A detailed derivation of the algorithm in its current form and more discussions can be found in [sun2024accelerating] and the references therein.
@@
Note that Step 4 in Algorithm 2 is from the Halpern iteration with a stepsize of $\frac{1}{k+2}$. Without Step 4, the HOT algorithm reduces to the ADMM with a unit step size. According to [sun2024accelerating, Corollary 3.5], we can obtain the global convergence of the HOT algorithm in the following proposition. The proof can be found in Appendix B.
@@
for any $w=(y,z,x)\in\mathbb{R}^{M_{3}}\times\mathbb{R}^{N}\times\mathbb{R}^{N}$. Note that $\mathcal{R}(w^{*})=0$ is equivalent to the facts that $x^{*}\in\mathbb{R}^{N}$ and $(y^{*},z^{*})\in\mathbb{R}^{M_{3}}\times\mathbb{R}^{N}$ are the solution to problems (8) and (9), respectively. The KKT residual $\|\mathcal{R}(\cdot)\|$ is a commonly used and practical measure for the quality of the approximation solution to (8). It follows from [sun2024accelerating, Theorem 3.7] that the HOT algorithm enjoys an appealing $O(1/k)$ nonergodic convergence rate in terms of the KKT residual for solving (8), which is summarized in the following proposition. The proof can be found in Appendix C.
@@
$\|\mathcal{R}(\bar{w}^{k})\|\leq\left(\frac{\sigma+1}{{\sigma}}\right)\frac{R_{0}}{(k+1)}$ (14)
and
$-\|z^{*}\|\frac{R_{0}}{(k+1)}\leq\langle c,{\bar{x}}^{k}-x^{*}\rangle\leq(\sigma\|z^{*}\|+R_{0})\frac{R_{0}}{\sigma(k+1)}.$
@@
Note that, without acceleration, the ADMM has an $O(1/\sqrt{k})$ non-ergodic rate in terms of both the objective function value gap and feasibility violations [davis2016convergence, cui2016convergence]. In contrast, the HOT algorithm in Algorithm 2 achieves an $O(1/k)$ non-ergodic convergence rate, offering significant advantages for solving large-scale OT problems.
@@ III-B A fast implementation of the HOT algorithm
Therefore, the main computational bottleneck of the HOT algorithm for solving (8) is solving the linear system (15). Note that the dimension of the matrix $AA^{\top}$ is $M_{3}\times M_{3}$ with $M_{3}$ defined in (4). In applications, $M_{3}$ is usually a huge number. As an illustrative example, for an image with $256\times 256$ pixels, $M_{3}=196,607$. As a result, it is not computationally affordable for computing a (sparse) Cholesky decomposition for the matrix $AA^{\top}$ or solving the linear system (15) with standard direct solvers. Indeed, it is computationally expensive even for computing the matrix $AA^{\top}$. Instead, in the remaining part of this subsection, we will derive a linear time complexity procedure for solving the linear equation $AA^{\top}{y}=R$ with a given vector $R\in\mathbb{R}^{M_{3}}$. It is worthwhile mentioning that our procedure does not require calculating nor storing the matrix $AA^{\top}$. Note that ${A}A^{\top}$ can be written in the following form:
@@
Remark 3.
Leveraging the sparsity of $A$, we can compute $Ax$ and $A^{\top}y$ with linear space complexity $O(N)$. Furthermore, according to Proposition 4, the linear system $AA^{\top}y=R$ can be solved with a memory cost of $O(M_{3})$. Hence, Algorithm 2 can be implemented with linear space complexity $O(N)$. When $m=n$, this corresponds to $O(M^{1.5})$.
@@
$O\left(\left(\frac{1+\sigma}{\sigma}\left(\left\|x^{0}-x^{*}\right\|+\sigma\left\|z^{0}-z^{*}\right\|\right)\right)\frac{m^{2}n+mn^{2}}{\varepsilon}\right).$
@@
Note that the complexity of the Sinkhorn method for solving the original OT problem (1) is $O\left(\frac{m^{2}n^{2}}{\varepsilon^{2}}\right)$ to achieve an $\varepsilon$-accuracy solution in terms of objective function value [dvurechensky2018computational]. Even by leveraging the separable structure of the cost function $c$ defined in (2), the convolutional Sinkhorn method, as mentioned in [solomon2015convolutional] and [auricchio2018computing], still requires $O\left(\frac{m^{2}n+mn^{2}}{\varepsilon^{2}}+m^{2}n^{2}\right)$ to compute the KW distance. In contrast, the HOT algorithm exhibits lower overall computational complexity, offering a significant advantage in calculating the KW distance between large-scale histograms.
Remark 5.
Assume that $m=n$. Blanchet et al. [blanchet2024towards] demonstrated that a running time of $\widetilde{O}({M^{2}}/{\varepsilon})$ is a bottleneck in their approach by reducing an instance of the bipartite matching problem (BMP) to an OT problem. While HOT achieves a complexity of $O(M^{1.5}/\varepsilon)$, it cannot be directly applied to solve the maximum cardinality BMP. This is because, in the reduction from BMP to the OT problem, the ground distance used in [blanchet2024towards] differs from the one in our paper. Consequently, the hardness result in [blanchet2024towards] does not directly contradict our complexity bound.
@@ IV Experiments
In this section, we comprehensively compare the HOT algorithm with five other state-of-the-art methods on the popular DOTmark dataset [schrieber2016dotmark]. Additionally, employing the transport plan derived from Algorithm 1, we test the performance of the reduced OT model on the color transfer task. The numerical results reveal that HOT provides significant advantages over state-of-the-art methods in both memory and computational efficiency, particularly for large-scale problems. More numerical results to compare the HOT algorithm to the W2NeuralDual method implemented in the OTT-JAX library [cuturi2022optimal] can be found in Appendix E. We also present an additional application of HOT for domain adaptation in Appendix F.
@@ IV-A Numerical comparison on the DOTmark dataset
The DOTmark dataset [schrieber2016dotmark] is a comprehensive collection of benchmark instances for evaluating and comparing algorithms in the field of optimal transport. It consists of a variety of instances categorized into different classes, such as Classic Images, Shapes, and Gaussian Distributions. In this experiment, we selected eight images each from the Classic Images and Shapes categories, as illustrated in Fig. 2. These selected images were resized to four different resolutions: $64\times 64$, $128\times 128$, $256\times 256$, and $512\times 512$. Finally, we randomly selected 10 pairs from each category and computed the KW distance for each pair.
@@
  • Sinkhorn [cuturi2013sinkhorn] is a widely used algorithm for computing an approximate KW distance by solving an entropy-regularized OT problem. As the Introduction Section discusses, its performance is sensitive to the regularization parameter, denoted as $\lambda$. To achieve varying levels of solution accuracy, we selected the $\lambda$ to be $1\%$, $0.1\%$, and $0.01\%$ of the median transport cost, following the setup from [solomon2015convolutional]. Due to the potential numerical instability caused by small $\lambda$ values, we employed the Log-domain Sinkhorn algorithm implemented by POT [flamary2021pot] as the baseline method.
  • Convolutional Sinkhorn [solomon2015convolutional, solomon2018optimal] is an improved version of the Sinkhorn algorithm, optimized for computing distances over regular two-dimensional grids. Specifically, since the kernel matrix used by the Sinkhorn algorithm can be constructed using a Kronecker product of two smaller matrices, a matrix-vector product using a matrix of dimension $M\times M$ can be replaced by two matrix-matrix products over matrices of dimension $\sqrt{M}\times\sqrt{M}$, resulting in a significant improvement in computational efficiency. In our experiment, based on the MATLAB implementation of the convolutional Sinkhorn in [auricchio2018computing], we further developed the Log-domain convolutional Sinkhorn using PyTorch. The regularization parameter $\lambda$ is kept the same as that of the original Sinkhorn method.
@@
  • Network Simplex implemented in the Lemon C++ graph library  [^3] https://lemon.cs.elte.hu/, is a highly efficient algorithm for solving uncapacitated minimum cost flow problems [kovacs2015minimum]. It has demonstrated the computational advantages over the Sinkhorn type methods in [auricchio2018computing] for small size images by solving the reduced model (8).
  • ADMM [gabay1976dual, glowinski1975approximation] is a popular first-order primal-dual method for solving large-scale optimization problems. It has shown great potential in solving large-scale optimal transport problems (1) in a GPU setting [mai2022fast]. In this numerical experiment, we use a generalized ADMM [eckstein1992douglas, xiao2018generalized] to solve the equivalent model (8) (replacing Step 4 in Algorithm 2 with $w^{k+1}=(1-\rho)w^{k}+\rho{\bar{w}}^{k}$ and setting $\rho=1.7$) as the baseline to evaluate the acceleration effect of the Halpern iteration.
@@
$\small{\rm KKT_{res}}=\max\left\{\frac{\|A^{\top}y+z-c\|}{1+\|c\|},\frac{\|\min(x,z)\|}{1+\|x\|+\|z\|},\frac{\|Ax-b\|}{1+\|b\|}\right\}.$ (29)
@@
We present the average results of 10 pairs within each category for all tested algorithms in Table II. Since Gurobi runs out of memory for images sized from $256\times 256$ to $512\times 512$, we only report the feaserr for the HOT, Network Simplex, and ADMM. Additionally, Table II only shows the results of Sinkhorn-type methods with $\lambda=0.01\%$ of the median transport cost, as this parameter returns a solution of comparable quality to other methods. For more results of the convolutional Sinkhorn with different regularization parameters, refer to Table III. Due to space constraints, we omit the results of the Sinkhorn with varying parameters, which exhibit similar performance to the convolutional Sinkhorn in terms of the gap.
@@
    Computational efficiency: HOT can return a comparable solution in terms of the gap and feaserr in the shortest time. Although the Network Simplex and Gurobi have computational advantages for computing the KW distance for small-scale images, these methods cannot handle large-scale problems effectively. IPM implemented in Gurobi requires solving a linear equation that relies heavily on Cholesky decomposition, causing the computational cost of each iteration to increase rapidly with image size. Additionally, the inherent sequential nature of the Network Simplex algorithm makes it challenging to parallelize effectively. In contrast, HOT and convolutional Sinkhorn benefit from low per-iteration costs and are easily parallelizable. However, the convolutional Sinkhorn needs to recover the solution to the original problem (1) to compute the KW distance, making it unsuitable for large-scale problems. As a result, for images sized $128\times 128$, HOT achieves a 17.44x speedup over Network Simplex, a 15.83x speedup over Gurobi, and a 19.54x speedup over convolutional Sinkhorn. Additionally, compared to ADMM, HOT benefits from a superior $O(1/k)$ iteration complexity and saves 40% of iterations for images sized $512\times 512$.
@@
Category Resolution HOT Network Simplex Gurobi ADMM Convolutional Sinkhorn Sinkhorn
@@
TABLE III: The numerical results of the convolutional Sinkhorn method with different $\lambda$.
@@
Convolutional Sinkhorn Classic $64\times 64$ time(s) 16.18 1.60 0.19
@@
To further illustrate the benefit of the explicit solution of the linear system presented in Proposition 4, we conduct a comparison between solving the linear equation using Proposition 4 and sparse Cholesky decomposition. Given that the average iteration number of HOT is around 1500, we solve the linear system (17) 1500 times for different image sizes. The results are shown in Fig. 3. It is clear that the computational costs of Cholesky decomposition increase rapidly as image size increases. Additionally, the forward-backward substitution method used to solve the linear system is also time-consuming because it cannot be efficiently parallelized. In contrast, solving linear systems using Proposition 4 can be easily parallelized. It consumes significantly less time and remains constant as the size varies because it only requires $O(M_{3})$ flops, as shown in Corollary 1. Additional numerical results on transport plan recovery time, GPU vs. CPU comparisons, and the acceleration effects of Halpern iteration for high-accuracy solutions can be found in Appendix D.